# Machine Learning and Feature Ranking for Impact Fall Detection Event Using Multisensor Data

### v1

@@ I Introduction
According to the World Health Organization report, falls affect 32 percent of older adults annually, making it a leading cause of mortality among this demographic [1]. The elderly population is particularly vulnerable to the detrimental effects of falls, which can result in fractures and cognitive impairments[2]. Recent research efforts have focused on fall detection to predict and prevent such incidents [3]. The primary objective of these studies is to minimize fall-related injuries and provide timely assistance to individuals at risk. An important advancement in this field is the implementation of pre-impact fall detection systems[4]. These systems identify falls at an early stage before the impact occurs. To detect falls across different stages, including pre-impact, and post-impact, accelerometer-based fall detection systems have gained prominence. However, while identifying fall at the early stage and post-stage can help to rescue people, it is crucial to identify the impact point, in order to avoid false alert and accurately provide assistance to people who really fall. Furthermore, it has been observed that the current datasets available for fall detection lack proper preprocessing, leading to inaccurate detection of falls and impacts in real-life scenarios. Therefore, depending on raw data from these datasets for the development of fall detection systems can lead to an increased occurrence of false positives or false negatives, as illustrated in Fig. 1(b). This figure demonstrates the occurrence of a false positive, where the system incorrectly identifies an impact where it has not actually occurred. This can cause false alarms and that reduces trust in the detection system. False negative cases are more dangerous in this context, where the person has actually fallen but the system does not detect it. Moreover, the detection of impacts from these datasets may suffer from delays, which further undermines the reliability of fall detection systems as shown in Fig. 1(a).
@@ III-D IMPACT DETECTION USING MACHINE LEARNING
SVM
{’C’: 1.0, ’kernel’: ’rbf’, ’gamma’: ’scale’}
Logistic Regression
{’C’: 1.0, ’solver’: ’lbfgs’}
Decision Tree
{’criterion’: ’gini’, ’max_depth’: None}
K-Nearest Neighbors
{’n_neighbors’: 5, ’weights’: ’uniform’}
Gaussian Naive Bayes
Random Forest
{’n_estimators’: 100, ’criterion’: ’gini’}
Stochastic Gradient Descent
{’alpha’: 0.0001, ’max_iter’: 1000}
Gradient Boosting
{’n_estimators’: 100, ’learning_rate’: 0.1}
@@
In the context of fall and impact detection,
@@
The SGD[20] algorithm learns to distinguish between the two classes (impacts and non-impacts) by iteratively adjusting the classifier’s parameters based on the input features provided.
@@ IV EXPERIMENTAL RESULTS
Algorithm Accuracy (%) Recall (%) Precision (%) F1-Score (%)
Training Time (s)
SVM 99.50 99.50 99.50 99.50
0.059
RF 99.28 98.47 99.18 98.47
0.607
SGD
94.47
97.71
95.85
96.77
0.013
NB
98.85
97.57
97.94
97.76
0.001
DT
98.85
97.57
97.94
97.76
0.034
KNN
98.35
98.45
99.03
98.74
0.005
LR
95.75
87.71
91.20
94.64
0.010
GBOOST 99.35 98.65 98.84 98.74
0.844
@@ V CONCLUSION
In conclusion, this study highlights the significance of appropriate pre-processing techniques in enhancing the accuracy and reliability of impact detection in fall events. The evaluation of various machine learning models revealed SVM as the most promising model for impact detection due to its high accuracy, robustness, and shorter inference time. These qualities make SVM suitable for fall impact identification in real-time. The findings of this study contribute to the advancement of fall detection technologies by highlighting the significance of pre-processing techniques and identifying effective machine learning models for impact detection. Moving forward, future work should focus on exploring the integration of multimodality by combining data from different sources. Additionally, we aim to incorporate deep learning algorithms to stay up-to-date with the latest advancements in impact detection algorithms. Moreover, the proposed pre-processing technique should be applied to new datasets, including real-world data, to thoroughly assess its performance and further enhance the accuracy, reliability, and adaptability of impact and fall detection systems.

### v2

@@ I Introduction
According to the World Health Organization report, falls affect 32 percent of older adults annually, making it a leading cause of mortality among this demographic [1]. The elderly population is particularly vulnerable to the detrimental effects of falls, which can result in fractures and cognitive impairments[2]. Recent research efforts have focused on fall detection to predict and prevent such incidents [3]. The primary objective of these studies is to minimize fall-related injuries and provide timely assistance to individuals at risk. An important advancement in this field is the implementation of pre-impact fall detection systems[4]. These systems identify falls at an early stage before the impact occurs.
Advancements in computer vision, driven by graphs, statistical techniques, and deep learning, have greatly enhanced visual data processing, which is particularly beneficial for improving fall detection accuracy [5, 6, 7, 8]. To detect falls across different stages, including pre-impact, and post-impact, accelerometer-based fall detection systems have gained prominence. However, while identifying fall at the early stage and post-stage can help to rescue people, it is crucial to identify the impact point, in order to avoid false alert and accurately provide assistance to people who really fall. Furthermore, it has been observed that the current datasets available for fall detection lack proper preprocessing, leading to inaccurate detection of falls and impacts in real-life scenarios. Therefore, depending on raw data from these datasets for the development of fall detection systems can lead to an increased occurrence of false positives or false negatives, as illustrated in Fig. 1(b). This figure demonstrates the occurrence of a false positive, where the system incorrectly identifies an impact where it has not actually occurred. This can cause false alarms and that reduces trust in the detection system. False negative cases are more dangerous in this context, where the person has actually fallen but the system does not detect it. Moreover, the detection of impacts from these datasets may suffer from delays, which further undermines the reliability of fall detection systems as shown in Fig. 1(a).
@@ III-D IMPACT DETECTION USING MACHINE LEARNING
SVM {’C’: 1.0, ’kernel’: ’rbf’, ’gamma’: ’scale’}
Logistic Regression {’C’: 1.0, ’solver’: ’lbfgs’}
Decision Tree {’criterion’: ’gini’, ’max_depth’: None}
K-Nearest Neighbors {’n_neighbors’: 5, ’weights’: ’uniform’}
Gaussian Naive Bayes -
Random Forest {’n_estimators’: 100, ’criterion’: ’gini’}
Stochastic Gradient Descent {’alpha’: 0.0001, ’max_iter’: 1000}
Gradient Boosting {’n_estimators’: 100, ’learning_rate’: 0.1}
@@
In the context of fall and impact detection, GB[23] can be a suitable algorithm due to its ability to handle sequential learning and refine the impact detection process over iterations.
  • SGD Classifier
@@ IV EXPERIMENTAL RESULTS
Algorithm Accuracy (%) Recall (%) Precision (%) F1-Score (%) Training Time (s)
SVM 99.50 99.50 99.50 99.50 0.059
RF 99.28 98.47 99.18 98.47 0.607
SGD 94.47 97.71 95.85 96.77 0.013
NB 98.85 97.57 97.94 97.76 0.001
DT 98.85 97.57 97.94 97.76 0.034
KNN 98.35 98.45 99.03 98.74 0.005
LR 95.75 87.71 91.20 94.64 0.010
GBOOST 99.35 98.65 98.84 98.74 0.844
@@ V CONCLUSION
In conclusion, this study highlights the significance of appropriate pre-processing techniques in enhancing the accuracy and reliability of impact detection in fall events. The evaluation of various machine learning models revealed SVM as the most promising model for impact detection due to its high accuracy, robustness, and shorter inference time. These qualities make SVM suitable for fall impact identification in real-time or time-sensitive scenarios. The findings of this study contribute to the advancement of fall detection technologies by highlighting the significance of pre-processing techniques and identifying effective machine learning models for impact detection. Moving forward, future work should focus on exploring the integration of multimodality by combining data from different sources. Additionally, the proposed pre-processing technique should be applied to new datasets to assess its performance and improve the accuracy, reliability, and adaptability of impact and fall detection systems.